OpenStudy (anonymous):

The sum of -4 and 8 times a number is multiplied by 3. The result is 15 less than 25 times the number. What is the number? a. -12 b. 6/7 c. 0 d. 3 e. none of the above

OpenStudy (shadowlegendx.):

This is simply changing words into an equation/expression

OpenStudy (anonymous):

i got none of the above.

OpenStudy (shadowlegendx.):

Do you know what the "sum" means?

OpenStudy (anonymous):

yes. sum is the answer.

OpenStudy (anonymous):

the total.

OpenStudy (texaschic101):

the sum means add

OpenStudy (anonymous):

yeah i know that. its the total in an addition problem though

OpenStudy (shadowlegendx.):

\[-4 + 8 \times y \times 3 = 25 - 15 \times y\]

OpenStudy (anonymous):

That's the exact same problem i wrote out from the question.

OpenStudy (shadowlegendx.):

Ok good!

OpenStudy (anonymous):

but when i plugged in the numbers i didn't get anything that matched

OpenStudy (shadowlegendx.):

Now we just do PEMDAS and like terms to solve

OpenStudy (shadowlegendx.):

-4 + 24y = 10y

OpenStudy (shadowlegendx.):

From that we would just further isolate the variable to get \[y = \frac{ 2 }{ 7 }\]

OpenStudy (anonymous):

that's what i got as well, which isn't one of the answer choices. so that would make E. none of the above correct, right?

OpenStudy (shadowlegendx.):

As we can all see, that is not one of the answers so it must fall into the "none of the above" category :)

OpenStudy (shadowlegendx.):

Yes, you are correct!

OpenStudy (anonymous):

alright, i had all the right steps then. i'm just trying to check over my answers because i'm studying for an upcoming exam right now. thank you :)